OpenStudy (anonymous):

f(x)=x^(1/x) how do i find the max value using calculus?

OpenStudy (anonymous):

1. We say that f(x) has an absolute (or global) maximum at if for every x in the domain we are working on. 2. We say that f(x) has a relative (or local) maximum at if for every x in some open interval around . 3. We say that f(x) has an absolute (or global) minimum at if for every x in the domain we are working on. 4. We say that f(x) has a relative (or local) minimum at if for every x in some open interval around .

OpenStudy (anonymous):

its really easy hope that helps

OpenStudy (anonymous):

no not really, i need to find out how to get a number out of it.

OpenStudy (anonymous):

@ArmyBoy97 you can't just copy and paste and not know what it means.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

1/ take derivative 2/ let it =0 to find x 3/ plug back to the equation to find min/max

OpenStudy (anonymous):

thanks let me try it

OpenStudy (anonymous):

is the derivative 1/x(x)(lnx)-0

OpenStudy (anonymous):

=0 i mean

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

Let y = x^(1/x) ln both sides lny = 1/x lnx derivative both sides \[\dfrac{y'}{y}= (\dfrac{1}{x})'lnx+(lnx)'\dfrac{1}{x}\\y'=-y(\dfrac{lnx}{x^2}+\dfrac{1}{x^2})\] replace y to have \(y'=-x^{1/x}(lnx+1)/x^2\) this expression =0 iff lnx+1 =0 --> lnx =-1 --> x = 1/e that is the value of x which make the graph min or max. However, you have just 1 value of x, you can't define whether it is min or max You must take second derivative to consider whether the concave is up or down, if it is up--> this point is min; if it is down, --> this point is max. I mean:

OpenStudy (anonymous):

\[y=(x ^{1/2}(1-\ln(x))/x^2\]

OpenStudy (anonymous):

why x^1/2?? x^(1/x)

OpenStudy (anonymous):

oh oops k let me try again

OpenStudy (anonymous):

is it supposed to be y(\[y(1/x ^{2}-lnx/x ^{2})\]

OpenStudy (anonymous):

Oh yeah, you are right. My bad. hihihi...

OpenStudy (anonymous):

coool i was like damn what did i do wrong

OpenStudy (anonymous):

can you show me how to do the second derivative to

OpenStudy (anonymous):

From y'= y(1/x^2-lnx/x^2) y" = y'(1/x^2-lnx/x^2) + y(1/x^2-lnx/x^2)' replace y' and y after take derivative the last term